I previously entered a bad email address.
How do I solve the following equation for P?
P - (PD) = N

Answer by nyc_function(2626)

P - (PD) = N
Factor P on the left side.
P(1 - D) = N
Now divide both side by the quantity (1 - D).
P = N/(1 - D)